# Approximation and decomposition of attractors of a Hopfield neural network system

### v1

@@ Abstract
In this paper, the Parameter Switching (PS) algorithm is used to approximate numerically attractors of a Hopfield Neural Network (HNN) system. The PS algorithm is a convergent scheme designed for approximating attractors of an autonomous nonlinear system, depending linearly on a real parameter. Aided by the PS algorithm, it is shown that every attractor of the HNN system can be expressed as a convex combination of other attractors. The HNN system can easily be written in the form of a linear parameter dependence system, to which the PS algorithm can be applied. This work suggests the possibility to use the PS algorithm as a control-like or anticontrol-like method for chaos.
@@ 1 Introduction
with $n=3$, $x_{0}\in\mathbb{R}^{3}$, $a=10$ and $c=8/3$, can be expressed in the form (1)-(2), if one considers $p:=\rho$ as the bifurcation parameter, where
@@
As known, the global attractor (‘global minimal B-attractor’, ‘global uniform attractor’ or ‘maximal attractor’), represents a major research topic in dynamical systems, especially within the context of PDEs (see e.g. [10]). From the definition, a global attractor is a compact and invariant set containing all the dynamics evolving from all possible initial conditions. In other words, it contains all solutions, including stationary solutions, periodic solutions, as well as chaotic attractors, relevant to the asymptotic behaviors of the system. The term local attractor is assigned sometimes to attractors that are not global attractors.
@@
For a given parameter $p$ which is a function on initial condition $x_{0}$, the attractor may contain several local attractors, which only attract trajectories from a subset of initial conditions, specified by its basin of attraction (see e.g. [12]).
Neural networks, successors of analog computers, are useful in exploring the relationship between electrical activities and information transfer in the human brain to reveal the mechanisms of memory, association, and emotional expression [20, 17]. Neural networks are systems that typically are used to recognize, describe, or even to control the dynamics.
A chaotic system modeled by (1)-(2) is the classical Hopfield Neural Network (HNN) model, a highly significant model since it can emulate complex dynamic behaviors, similar to how the brain processes information [32] In 1980s, Hopfield proposed the HNN in the pioneering works [15][16], which combines statistical mechanics and can be derived from an electronic circuit. Fractional-order HNNs are considered in [24] and [5].
The HNN considered in this paper is a simplified popular variant of the simplest continuous-time model of a recurrent neural network that “realizes everything” in the sense that its dynamics can be determined with an arbitrary accuracy using its parameters [1], and presents a“maximal dynamic complexity” [34], which is modeled by the following Kirchhoff equations [16, 35] (see also [7, 23, 36]):
@@
$\begin{array}[l]{l}B=\begin{pmatrix}1&0&0\\ 0&0&0\\ 0&0&0\end{pmatrix}\\ \end{array}g(x)=W\begin{pmatrix}\tanh x_{1}\\ \tanh x_{2}\\ \tanh x_{3}\end{pmatrix}=\begin{pmatrix}2\tanh x_{1}-1.2\tanh x_{2}\\ -x_{2}+2\tanh x_{1}+1.71\tanh x_{2}+1.15\tanh x_{3}\\ -x_{3}+4.75\tanh x_{1}+1.1\tanh x_{3}\end{pmatrix}.$
@@
There are applications where transient chaos can be quite disastrous, such as the situations of species extinction or power-grid voltage collapse. As known, the notion of attractors as limit sets, as time goes to infinity. Numerically this characteristic is specified as a sufficiently large time integration. Therefore, if after a relatively long but reasonable integration time, chaos vanishes whereas the underlying transient should be called “set”, not “attractor”, which is an asymptotical approach. As shown in this paper and in [7], depending on the coupling coefficients $W$, the HNN system might have non-chaotic attractors but finite time chaotic transients.
Sometimes, the conversion of the transient chaos into sustained chaos can avoid catastrophes related to sudden chaos collapses (chaos anticontrol) [9]. However, there are situations where chaotic transients are undesirable, for which chaos control techniques are useful. A recently found phenomenon of transient chaos, doubly transient chaos, appears in many systems (chemical reactions, binary star behavior, etc.) and is likely far less predictable than previously thought and is fundamentally different from hyperbolic and nonhyperbolic transient chaos [27].
@@ 2 Parameter Switching Algorithm
In this section, the PS algorithm is briefly introduced (details can be found from the indicated references).
Throughout, the following reasonable mild conditions for most chaotic systems are imposed:
(H) Existence, uniqueness and boundness of solutions are assumed, and there exist only hyperbolic equilibria.
Note that, instead of the boundness, the dissipative property, an essential prerequisite for the existence of global attractors, can be used (see e.g. [2, 31]). Therefore, if $f\in C^{1}$, the dissipativity is ensured via the divergence (Gauss) theorem (see, Chapter 16.9 [30]).
Let $\mathcal{A}$ be the set of all attractors depending on parameter $p$ and $\mathcal{P\subset}\mathbb{R}$ be the set of the corresponding admissible values of $p$. Because of the solution uniqueness, for a chosen $x_{0}$ to each $p\in\mathcal{P}$, there corresponds a unique attractor $A_{p}\in\mathcal{A}$. Let $\mathcal{P}_{N}=\{p_{1},p_{2},\ldots,p_{N}\}\subset\mathcal{P}$, $p_{1}<p_{2}<\ldots,p_{N}$, a finite ordered set of $N\geq 2$ parameters, to which there corresponds the attractors set $\mathcal{A}_{N}=\{A_{p_{1}},A_{p_{2}},\ldots,A_{p_{N}}\}\subset{\mathcal{A}}$.
Because the assumption H, $\mathcal{A}$ is a non-empty set (see e.g. [22]). Therefore, for the considered class of systems, a bijection between the sets $\mathcal{P}$ and $\mathcal{A}$ can be naturally used. In this way, for any $p\in\mathcal{P}$, a unique attractor is defined and reversely.
The PS algorithm is a time-varying, periodically switching the parameter $p$ within a chosen set $\mathcal{P}_{N}$, according to some rule while the IVP is integrated with a convergent scheme for ODEs with fixed step-size $h$, on nodes $nh$, $n=1,2,\ldots$. Symbolically, the algorithm is expressed as the following scheme
@@
The scheme $S$ should be read as follows: while the attractor is numerically generated along the trajectory, the first $m_{1}$ integration steps of size $h$, i.e. for the $1st,2nd,\ldots,m_{1}th$ steps, the parameter $p$ is set $p_{1}$. Then, for the next $m_{2}$ steps, $(m_{1}+1)th,(m_{1}+2)th,\ldots(m_{1}+m_{2})th$, $p$ receives the value $p_{2}$ and the integration continues until $p=p_{N}$ for the steps $(m_{N-1}+1)th,(m_{N-1}+2)th,\ldots,(m_{N-1}+m_{N})th$. Next, the algorithm repeats on the further set of $N$ subintervals, and so on. Thus, the time periodicity of $p$ switching is $T=\sum_{i=1}^{N}m_{i}h$.
@@
    The selection of the sets $\mathcal{P}_{N}$ and weights $\mathcal{M}_{N}$ to obtain a desired value $p^{\circ}$ is not unique;
@@
The measure of the “success” of the attractor approximation, proved analytically by Theorem 3, can be verified numerically too. However, this task is usually a difficult one, because the size and the shape of an attractor might change once with the control parameter. Also, attractors geometric structure in $\mathbb{R}^{n}$ can be complicated. Considering the difficulties to compare attractors, the following criterion, which is more practical rather than rigorous theoretical, and which is a suitable modification and adaptation of the concept of topological equivalence [14] is introduced (see also [3], [11], [14], [18], [22]):
@@
An attractor $A_{p_{i}}$ is considered to approximate numerically the attractor $A_{p_{j}}$, for $i\neq j$, if their geometrical forms in the phase space (almost) coincide and the sense of the motion along the attractors is preserved.
@@
Given an attractor $A_{p^{\diamond}}$ there exist a set $\mathcal{P}_{N}$ and a set of weights $\mathcal{M}_{N}$, $N\geq 2$, such that $p^{\diamond}$ can be obtained from (6). Moreover, $A_{p^{\diamond}}$ is approximated by the switched attractor $A^{*}$ obtained with the PS algorithm.
@@
The numerical implementation of the PS algorithm has been realized via the fourth-order Runge-Kutta method (RK4) with the integration step $h$.
@@
    Consider the case of the Lorenz system and the set $\mathcal{P}_{4}=\{86,95,96,99\}$ with weights $\mathcal{M}_{4}=\{2,1,2,1\}$. Then, from Corollary 5, using scheme (5), $S=[m_{1}\cdot p_{1},m_{2}\cdot p_{2},m_{3}\cdot p_{3},m_{4}\cdot p_{4}]=[2\cdot 86,1\cdot 95,2\cdot 96,1\cdot 99]$, from (6) one obtains $p^{\circ}=(m_{1}\times p_{1}+m_{2}\times p_{2}+m_{3}\times p_{3}+m_{4}\times p_{4})/(m_{1}+m_{2}+m_{3}+m_{4})=93$, with the underlying averaged attractor $A_{p^{\circ}}:=A_{93}$ (stable cycle, blue plot in Fig. 1). The scheme $S$ reads as follows: for the first 2 integration steps, $p$ is set to $86$, the next step, $p=95$, and the next 2 steps, $p=96$, and so on, until the integration interval $I$ is covered. By applying the PS algorithm one obtains the switched attractor $A^{*}$ (red plot in Fig. 1). Transients of both attractors have been removed. As can be seen, there exists a perfect match between the two attractors.
@@
    For a given set $\mathcal{P}_{N}$, there exist several $N$ weights $m_{i}$, which give $p^{\circ}$ (Remark 2 ii)). Moreover, a specified value $p^{\circ}$ can be obtained with several sets of $\mathcal{P}_{N}$ and weights $\mathcal{M}_{N}$ with $N\geq 2$. For example, from Corollary 6, the attractor $A_{93}$ ($A^{\diamond}$ in Corollary 6) can be approximated with the scheme $S=[1\cdot 92,1\cdot 94]$, or $S=[2\cdot 86,1\cdot 95,3\cdot 97]$, which both give via (6) $p^{\diamond}:=p^{\circ}=93$.
@@
    Not only stable cycles can be approximated with the PS algorithm, but chaotic attractors too. For example, by switching (alternating) the parameter $p$ within the two elements set $\mathcal{P}_{2}=\{100,133\}$, with weights $m_{1}=m_{2}=1$, i.e. within every integration step, $p$ is alternated between the values of $100$ and $133$, one obtains the switched attractor $A^{*}$ (Fig. 2), which approximates the average chaotic attractor $A_{p^{\circ}}$, with $p^{\circ}=(100+133)/(1+1)=116.5$.
The simple Matlab code presented in [6] allows to generate via the standard RK method, the switched $y_{n}$ and averaged $x_{n}$ solutions, for a given system, function on the sets $\mathcal{P}_{N}$ and $\mathcal{M}_{N}$, initial conditions $x_{0}$ and $y_{0}$, and the integration step size $h$.
@@ 3 Attractor decomposition
The average parameter value $p^{\circ}$ given by (6) is a convex combinantion of the elements of the set $\mathcal{P}_{N}$.
@@
Consider the Lorenz system in Example 1. The attractors generated by the set $\mathcal{P}_{4}$, $\mathcal{A}_{4}=\{A_{86},A_{95},A_{96},A_{99}\}$, are plotted in Fig. 1 (a). The switched attractor $A^{*}$ approximates the average attractor $A_{p^{\circ}}$ with $p^{\circ}=93$. With $A_{86}=H(86)$, $A_{95}=H(95),A_{96}=H(96),A_{99}=H(99)$, and $\alpha_{k}$, $k=1,2,3,4$, given by (7), $\alpha_{1}=m_{1}/\sum_{k=1}^{4}m_{k}=2/6=1/3,\alpha_{2}=1/6,\alpha_{3}=1/6$. From the relation (9), one obtains
@@
[6] Giving the ordered set $\mathcal{P}_{N}=\{p_{min},\ldots,p_{max}\}$, using scheme (6), $p^{\circ}$ satisfies the relation $p^{\circ}\in(p_{min},p_{max})$. Therefore, via $H$, the order induced by $\mathcal{P}_{N}$ in the set $\mathcal{A}_{N}=\{A_{min},\ldots,A_{max}\}$, implies that the switched attractor $A^{*}$, which approximates the average attractor $A_{p^{\circ}}$, verifies $A^{*}\in(A_{min},A_{max})$. Also, by the convexity of $p^{\circ}$, $A^{*}$ and $A_{p^{\circ}}$ are different from all attractors in $\mathcal{A}_{N}$.

### v2

@@ Abstract
In this paper, the Parameter Switching (PS) algorithm is used to numerically approximate attractors of a Hopfield Neural Network (HNN) system. The PS algorithm is a convergent scheme designed for approximating the attractors of an autonomous nonlinear system, depending linearly on a real parameter. Aided by the PS algorithm, it is shown that every attractor of the HNN system can be expressed as a convex combination of other attractors. The HNN system can easily be written in the form of a linear parameter dependence system, to which the PS algorithm can be applied. This work suggests the possibility to use the PS algorithm as a control-like or anticontrol-like method for chaos.
@@ 1 Introduction
with $n=3$, $x_{0}\in\mathbb{R}^{3}$, $\sigma=10$ and $\beta=8/3$, can be expressed in the form (1)-(2), if one considers $p:=\rho$ as the bifurcation parameter, where
@@
As is known, the global attractor (‘global minimal B-attractor’, ‘global uniform attractor’ or ‘maximal attractor’), represents a major research topic in dynamical systems, especially within the context of partial differential equations (see, e.g. [8]). From the definition, a global attractor is a compact and invariant set containing all the dynamics evolving from all possible initial conditions. In other words, it contains all solutions, including stationary solutions, periodic solutions, as well as chaotic attractors, relevant to the asymptotic behaviors of the system. The term local attractor is sometimes assigned to attractors that are not global attractors.
@@
For a given parameter $p$ which is a function of initial condition $x_{0}$, the attractor may contain several local attractors, which only attract trajectories from a subset of initial conditions, specified by its basin of attraction (see e.g. [9]).
Neural networks, the successors of analog computers, are useful in exploring the relationship between electrical activities and information transfer in the human brain to reveal the mechanisms of memory, association, and emotional expression [13, 14]. Neural networks are systems that are typically used to recognize, describe, or even to control the dynamics.
A chaotic system modeled by (1)-(2) is the classical Hopfield Neural Network (HNN) model, a highly significant model since it can emulate complex, dynamic behaviors, similar to how the brain processes information [15] In the 1980s, Hopfield proposed the HNN in his pioneering works [16][17], which combines statistical mechanics and can be derived from an electronic circuit. Fractional-order HNNs are considered in [18] and [4].
The HNN considered in this paper is a simplified popular variant of the simplest continuous-time model of a recurrent neural network that “realizes everything” [19] in the sense that its dynamics can be determined with an arbitrary accuracy using its parameters, and presents a“maximal dynamic complexity” [20], which is modeled by the following Kirchhoff equations [17, 21] (see also [5, 22, 23]):
@@
$\begin{array}[l]{l}B=\begin{pmatrix}1&0&0\\ 0&0&0\\ 0&0&0\end{pmatrix}\\ \end{array}g(x)=W\begin{pmatrix}\tanh x_{1}\\ \tanh x_{2}\\ \tanh x_{3}\end{pmatrix}=\begin{pmatrix}2\tanh x_{1}-1.2\tanh x_{2}\\ -x_{2}+1.9995\tanh x_{1}+1.71\tanh x_{2}+1.15\tanh x_{3}\\ -x_{3}-4.75\tanh x_{1}+1.1\tanh x_{3}\end{pmatrix}.$
@@
There are applications where transient chaos can be quite disastrous, such as the situation of species extinction or power-grid voltage collapse. As known, the notion of attractors as limit sets, as time goes to infinity. Numerically, this characteristic is specified as a sufficiently large time integration. Therefore, if after a relatively long but reasonable integration time, chaos vanishes, whereas the underlying transient should be called “set”, not “attractor”, which is an asymptotical approach. As shown in this paper and in [5], depending on the coupling coefficients $W$, the HNN system might have non-chaotic attractors but finite time chaotic transients.
Sometimes, the conversion of transient chaos into sustained chaos can avoid catastrophes related to sudden chaos collapses (chaos anticontrol) [24]. However, there are situations where chaotic transients are undesirable, for which chaos control techniques are useful. A recently found phenomenon of transient chaos, doubly transient chaos, term used to emphasize the fact that in many systems, chaotic solutions finally stop (chemical reactions, binary star behavior, etc.) and is likely far less predictable than previously thought and is fundamentally different from hyperbolic and nonhyperbolic transient chaos [25].
@@
The manuscript is organized as follows: Section 1, Introduction; Section 2, where the PS algorithm is briefly described; Section 3, where it is shown how attractors can be decomposed as convex combination of other attractors of the considered system; Section 4, where the PS algorithm is applied to approximate attractors of the considered HNN system and also, to decompose HNN’s attractors as convex combinations of other attractors and the last section, Conclusions and Discussion Section. In Appendix is presented the Matlab code, which allows to use randomly the PS in the case of the HNN system.
@@ 2 Parameter Switching Algorithm
In this section, the PS algorithm is briefly introduced (details can be found in the indicated references).
Throughout, the following reasonable, mild conditions for most chaotic systems are imposed:
(H) Existence, uniqueness, and boundness of solutions are assumed, and there exist only hyperbolic equilibria.
Note that, instead of boundness, the dissipative property, an essential prerequisite for the existence of global attractors, can be used (see, e.g. [28, 10]). Therefore, if $f\in C^{1}$, the dissipativity is ensured via the divergence (Gauss) theorem (see, Chapter 16.9 [29]).
Let $\mathcal{A}$ be the set of all attractors depending on parameter $p$ and $\mathcal{P\subset}\mathbb{R}$ be the set of the corresponding admissible values of $p$. Because of the solution uniqueness, for a chosen $x_{0}$ for each $p\in\mathcal{P}$, there corresponds a unique attractor $A_{p}\in\mathcal{A}$. Let $\mathcal{P}_{N}=\{p_{1},p_{2},\ldots,p_{N}\}\subset\mathcal{P}$, $p_{1}<p_{2}<\ldots,p_{N}$, a finite ordered set of $N\geq 2$ parameters, to which there corresponds the attractors set $\mathcal{A}_{N}=\{A_{p_{1}},A_{p_{2}},\ldots,A_{p_{N}}\}\subset{\mathcal{A}}$.
Because of the assumption H, $\mathcal{A}$ is a non-empty set (see, e.g. [30]). Therefore, for the considered class of systems, a bijection between the sets $\mathcal{P}$ and $\mathcal{A}$ can be naturally used. In this way, for any $p\in\mathcal{P}$, a unique attractor is defined and reversely.
The PS algorithm is a time-varying, periodically switching the parameter $p$ within a chosen set $\mathcal{P}_{N}$, according to some rules, while the IVP is integrated with a convergent scheme for ordinary differential equations with a fixed step-size $h$, on nodes $nh$, $n=1,2,\ldots$. Symbolically, the algorithm is expressed as the following scheme
@@
The scheme $S$ should be read as follows: while the attractor is numerically generated along the trajectory, the first $m_{1}$ integration steps of size $h$, i.e. for the $1st,2nd,\ldots,m_{1}th$ steps, the parameter $p$ is set $p_{1}$. Then, for the next $m_{2}$ steps, $(m_{1}+1)th,(m_{1}+2)th,\ldots(m_{1}+m_{2})th$, $p$ receives the value $p_{2}$ and the integration continues until $p=p_{N}$ for the steps $(m_{N-1}+1)th,(m_{N-1}+2)th,\ldots,(m_{N-1}+m_{N})th$. Next, the algorithm repeats on the next set of $N$ subintervals, and so on. Thus, the time periodicity of $p$ switching is $T=\sum_{i=1}^{N}m_{i}h$.
@@
    The selection of the sets $\mathcal{P}_{N}$ and weights $\mathcal{M}_{N}$ to obtaining a desired value $p^{\circ}$ is not unique;
@@
The measure of the “success” of the attractor approximation, proved analytically by Theorem 3, can be verified numerically too. However, this task is usually a difficult one, because the size and shape of an attractor might change once with the control parameter. Also, attractors geometric structure in $\mathbb{R}^{n}$ can be complicated. Considering the difficulties to comparing attractors, the following criterion, which is more practical rather than rigorous theoretical, and which is a suitable modification and adaptation of the concept of topological equivalence [32] is introduced (see also [33], [34], [32], [35], [30]):
@@
An attractor $A_{p_{i}}$ is considered to approximate numerically the attractor $A_{p_{j}}$, for $i\neq j$, if their geometrical forms in the phase space (almost) coincide and the sense of motion along the attractors is preserved.
@@
Given an attractor $A_{p^{\diamond}}$ there exists a set $\mathcal{P}_{N}$ and a set of weights $\mathcal{M}_{N}$, $N\geq 2$, such that $p^{\diamond}$ can be obtained from (6). Moreover, $A_{p^{\diamond}}$ is approximated by the switched attractor $A^{*}$ obtained with the PS algorithm.
@@
Even though one of the convergence proofs of the PS algorithm is not related to some numerical scheme [2], the explicit numerical approach of the proof in [3] allows the implementation of the algorithm via some fixed step-size scheme (in this paper, the Runge-Kutta method (RK4)) with the integration step $h$.
@@
    Consider the case of the Lorenz system and the set $\mathcal{P}_{4}=\{86,95,96,99\}$ with weights $\mathcal{M}_{4}=\{2,1,2,1\}$. Then, from Corollary 5, using scheme (5), $S=[m_{1}\cdot p_{1},m_{2}\cdot p_{2},m_{3}\cdot p_{3},m_{4}\cdot p_{4}]=[2\cdot 86,1\cdot 95,2\cdot 96,1\cdot 99]$, from (6) one obtains $p^{\circ}=(m_{1}\times p_{1}+m_{2}\times p_{2}+m_{3}\times p_{3}+m_{4}\times p_{4})/(m_{1}+m_{2}+m_{3}+m_{4})=93$, with the underlying averaged attractor $A_{p^{\circ}}:=A_{93}$ (stable cycle, blue plot in Fig. 1). The scheme $S$ reads as follows: For the first 2 integration steps, $p$ is set to $86$, the next step, $p=95$, and the next 2 steps, $p=96$, and so on, until the integration interval $I$ is covered. Applying the PS algorithm, one obtains the switched attractor $A^{*}$ (red plot in Fig. 1). Transients of both attractors have been removed. As can be seen, there is a perfect match between the two attractors.
@@
    For a given set $\mathcal{P}_{N}$, there exist several $N$ weights $m_{i}$, which give $p^{\circ}$ (Remark 2 ii)). Moreover, a specified value $p^{\circ}$ can be obtained with several sets of $\mathcal{P}_{N}$ and weights $\mathcal{M}_{N}$ with $N\geq 2$. For example, from Corollary 6, the attractor $A_{93}$ ($A_{p^{\diamond}}$ in Corollary 6) can be approximated with the scheme $S=[1\cdot 92,1\cdot 94]$, or $S=[2\cdot 86,1\cdot 95,3\cdot 97]$, which both give via (6) $p^{\diamond}:=p^{\circ}=93$.
@@
    Not only stable cycles can be approximated with the PS algorithm, but chaotic attractors too. For example, by switching (alternating) the parameter $p$ within the two elements set $\mathcal{P}_{2}=\{100,133\}$, with weights $m_{1}=m_{2}=1$, i.e., within every integration step, $p$ is alternated between the values of $100$ and $133$, one obtains the switched attractor $A^{*}$ (Fig. 2), which approximates the average chaotic attractor $A_{p^{\circ}}$, with $p^{\circ}=(100+133)/(1+1)=116.5$.
The simple Matlab code presented in [1] allows to generate via the standard RK method, the switched $y_{n}$ and averaged $x_{n}$ solutions, for a given system, function on the sets $\mathcal{P}_{N}$ and $\mathcal{M}_{N}$, initial conditions $x_{0}$ and $y_{0}$, and the integration step size $h$. In this paper, for the Lorenz system, the set of initial conditions $x_{0}$ and $y_{0}$ of the switched solution and average solution, are $(1,1,1)$ and $(1.01,1.01,1.01)$, respectively.
@@ 3 Attractor decomposition
The average parameter value $p^{\circ}$ given by (6) is a convex combination of the elements of the set $\mathcal{P}_{N}$.
@@
In [1] is proved that the operation $\oplus$ is commutative and associative. As a consequence, the order of the considered parameters in the scheme (5) is not important.
@@
Consider the Lorenz system in Example 1. The attractors generated by the set $\mathcal{P}_{4}$, $\mathcal{A}_{4}=\{A_{86},A_{95},A_{96},A_{99}\}$, are plotted in Fig. 1 (a). The switched attractor $A^{*}$ approximates the average attractor $A_{p^{\circ}}$ with $p^{\circ}=93$. With $A_{86}=H(86)$, $A_{95}=H(95),A_{96}=H(96),A_{99}=H(99)$, and $\alpha_{k}$, $k=1,2,3,4$, given by (7), $\alpha_{1}=m_{1}/\sum_{k=1}^{4}m_{k}=2/6=1/3,\alpha_{2}=1/6,\alpha_{3}=1/3$ and $\alpha_{4}=1/6$. From the relation (9), one obtains
@@
[1] Giving the ordered set $\mathcal{P}_{N}=\{p_{min},\ldots,p_{max}\}$, using scheme (6), $p^{\circ}$ remains within the parameter interval $p^{\circ}\in(p_{min},p_{max})$. Therefore, similarly, via $H$ the order induced by $\mathcal{P}_{N}$ in the set $\mathcal{A}_{N}=\{A_{min},\ldots,A_{max}\}$, implies that the switched attractor $A^{*}$, which approximates the average attractor $A_{p^{\circ}}$, verifies $A^{*}\in(A_{min},A_{max})$. Also, from the convexity of $p^{\circ}$, $A^{*}$ and $A_{p^{\circ}}$ are different from all attractors in $\mathcal{A}_{N}$.